OpenStudy (anonymous):

Someone please help me with this, I have no idea how to do it. If sin theta =7/8 for pi/2 < or = theta < or = pi and cos alpha= - 5/6 for pi < or = pi < or = alpha < or = 3pi/ 2 find the value of cos ( theta - alpha ) PLEASE HELP ME!

OpenStudy (mertsj):

I'm going to use x and y for theta and alpha.

OpenStudy (mertsj):

\[\cos (x-y)=\cos x \cos y+\sin x \sin y\]

OpenStudy (mertsj):

I see that theta (I'm using x) is in the second quadrant.